What is the wavelength of a 5MHz

ultrasound pulse in soft tissue?

Assume that the speed of ultrasound

in soft tissue is 1500m/s. [Best

answerl

A. 0.3 mm

B. 0.5 mm

C. 3 mm

D. 5 mm


1
Expert's answer
2021-12-03T12:29:40-0500
"\lambda=\frac{v}{f}\\\lambda=\frac{1500}{5\cdot10^6}=0.3\ mm"